# Remote Work and Women’s Labor Supply: The New Gender Division at Home

### v1

@@ D. Placebo: Moving Covid shock to March 2018
  • Notes: We use data from the the 2017-2024 ATUS panel datasets. We show results for the time spent in labor, household and personal activities available in ATUS. Time-use variables are recorded by one member in the household. Then, as Section 2 points out, we have two draws from ATUS: the men-respondent sample and the women-respondent sample. While the former has data on the time spent doing activities by husbands, the latter will have for wives. There is a first panel that has data on the time spent in labor production, capturing the share of the whole day - over 24 hours - that respondents spent: working, working at their main work, working at their main work from home and outside home, working at their secondary work and commuting to work. There is a panel on personal care (i.e. sleeping or taking a shower) and leisure (ie. going to religious events or sport activities). Lastly there is household panel with the time spent doing the following activities: housework, cooking, shopping, primary childcaring (report doing childcaring as main activity), childcaring while working, total childcaring as secondary activity while doing other activities, total childcaring (the sum of primary and secondary childcaring) and the share of secondary childcaring on total childcaring. We restrict our sample to weekdays and non-holiday days. All regressions include fixed effects for state-by-time and husband’s occupation. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level. We exploit a fake timing as in Table B.24.
Table B.35: Placebo Test: Effect of Husbands’ WFH Occupation Shocks on Childcare and Child Well-Being Outcomes (COVID Shock Moved to March 2018)
Childcare by Childcare by Childcare by Childcare by Childcare Paid for Log Childcare
Wife while Spouse while Grandparent Relative by Daycare Childcare Expenditure
Working Wife Works
$Exposed_{o(H)}\times Post_{t}$ 0.025 -0.012 -0.037* -0.046** -0.007 0.028 -0.006
(0.028) (0.030) (0.022) (0.022) (0.021) (0.027) (0.095)
Mean pre-treatment 0.319 0.487 0.422 0.462 0.255 0.478 4.636
Obs 44,531 44,531 63,829 63,829 63,829 50,889 24,551
$R^{2}$ 0.14 0.17 0.16 0.15 0.20 0.20 0.36
  • Notes: We replicate the analysis from Table 4 but we only use the 2017-2019 SIPP data, and artificially change the post-Covid definition to be after March 2018. All regressions include fixed effects for state-by-time, husband’s occupation, and husband’s industry. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level.
Table B.36: Effect of Husbands’ WFH Occupation Shocks on Child Well-Being and Activities
A. Entire sample
(0.086) (0.180) (0.234) (0.245) (0.059) (0.062) (0.060)
Mean pre-treatment 6.532 5.817 3.533 2.677 0.353 0.486 0.535
Obs 44,620 44,620 31,873 31,873 27,705 27,705 27,705
$R^{2}$ 0.18 0.23 0.25 0.24 0.27 0.32 0.30
Appendix C Discussion on Identification Strategy
Our empirical design estimates the causal effect of an increase in the husband’s work-from-home (WFH) feasibility on the wife’s labor market outcomes. However, in many datasets, including the CPS, occupational information is only recorded for currently employed individuals. This constraint introduces two primary limitations. First, we are unable to systematically observe the wife’s own occupation when she is not employed, preventing us from identifying her exposure to WFH exposure in non-employment spells. Second, this data limitation complicates the disentangling of the husband’s WFH-induced shock from any contemporaneous effects driven by changes in the wife’s own WFH exposure. While the latter concern can be partially addressed by restricting the sample to wives who are employed – allowing us to examine adjustments at the intensive margin – the former limitation precludes a comprehensive analysis of women’s own responses to shifts in their WFH exposure, as this measure is defined only for the employed subsample.
To understand the potential consequences of this limitation, consider a simplified version of our baseline specification:
$y_{i(H)t}=\alpha+\beta_{0}D^{o(H)}+\gamma D_{t}+\beta_{h}D^{o(H)}\times D_{t}+u_{i(H)t}$ (3)
@@ Appendix C Discussion on Identification Strategy
where $D^{o(H)}=1$ if the husband of woman $i$ is in a treated occupation (i.e., one with a large WFH shock), and $D_{t}=1$ for post-COVID observations.
@@
Suppose that unobserved variation in the wife’s own WFH exposure enters the error term:
$u_{i(H)t}=\beta_{1}D^{o(i)}+\beta_{w}D^{o(i)}\times D_{t}+\beta_{2}D^{o(i)}\times D^{o(H)}+\beta_{hw}D^{o(i)}\times D^{o(H)}\times D_{t}+\varepsilon_{i(H)t}$ (4)
where $D^{o(i)}=1$ if woman $i$ herself is in a treated occupation.
@@
Even if we assume that the cross-occupation interaction terms $\beta_{2}$ and $\beta_{hw}$ are zero, the difference-in-differences estimator $\hat{\beta}_{h}$ from  (3) still identifies:
[…]
Even under the assumption that assortative matching patterns are stable pre- and post-COVID (i.e., $\Delta P(D^{o(i)}=1|D^{o(H)}=1)=\Delta P(D^{o(i)}=1|D^{o(H)}=0)=0$), the estimator will still capture some direct effect through $\beta_{w}$.
While we cannot directly estimate or control for the wife’s WFH exposure across the full sample, we show in Section 5 that the observed effects are unlikely to be driven by changes in the wife’s occupational flexibility. For instance, we find that the husband’s WFH shock leads to changes in the wife’s intensive labor supply, whereas her own WFH exposure does not. This supports the interpretation of our estimates as capturing intra-household spillovers, rather than direct responses to the wife’s own occupational exposure.

### v2

@@ A. Entire sample
  • Notes: We use data from the the 2017-2024 ATUS panel datasets. We show results for the time spent in labor, household and personal activities available in ATUS. Time-use variables are recorded by one member in the household. Then, as Section 3 points out, we have two draws from ATUS: the men-respondent sample and the women-respondent sample. While the former has data on the time spent doing activities by husbands, the latter will have for wives. There is a first panel that has data on the time spent in labor production, capturing the share of the whole day - over 24 hours - that respondents spent: working, working at their main work, working at their main work from home and outside home, working at their secondary work and commuting to work. There is a panel on personal care (i.e. sleeping or taking a shower) and leisure (ie. going to religious events or sport activities). Lastly there is household panel with the time spent doing the following activities: housework, cooking, shopping, primary childcaring (report doing childcaring as main activity), childcaring while working, total childcaring as secondary activity while doing other activities, total childcaring (the sum of primary and secondary childcaring) and the share of secondary childcaring on total childcaring. We restrict our sample to weekdays and non-holiday days. All regressions include fixed effects for state-by-time and husband’s occupation. Controls include the wife’s age, education, and number of children, as well as the husband’s education. Standard errors are clustered at the husband’s occupation level. We exploit a fake timing as in Table B.31 by moving the arrival of the pandemic shock to March 2018.
Appendix C Discussion of Identification Strategy
Our empirical design estimates the causal effect of an increase in the husband’s WFH probability on the wife’s labor market outcomes. A key data limitation, however, is that in many surveys – including the ACS – occupational information is observed only for individuals currently employed. This constraint implies two challenges. First, we cannot systematically observe a wife’s own occupation when she is not employed, preventing us from measuring her WFH exposure during non-employment spells. Second, because women’s own exposure is unobserved for part of the sample, our baseline estimates necessarily combine the spillover effect arising from husbands’ WFH exposure with the direct effect operating through women’s own WFH exposure, which in the model is generated by assortative matching, as explained in Section 4.
To illustrate the implications, consider a simplified version of our baseline specification:
$y_{it}=\alpha+\beta_{0}D^{o(h(i))}+\gamma D_{t}+\beta_{h}D^{o(h(i))}\times D_{t}+u_{it},$ (8)
@@ Appendix C Discussion of Identification Strategy
where $D^{o(h(i))}=1$ if the husband of woman $i$ is in a treated occupation (i.e., one with a large WFH shock), and $D_{t}=1$ in post-COVID years.
@@
Suppose that unobserved variation in the wife’s own WFH exposure enters the error term as:
$u_{it}=\beta_{1}D^{o(i)}+\beta_{w}D^{o(i)}\times D_{t}+\beta_{2}D^{o(i)}\times D^{o(h(i))}+\beta_{hw}D^{o(i)}\times D^{o(h(i))}\times D_{t}+\varepsilon_{it},$ (9)
where $D^{o(i)}=1$ if the wife herself works in a treated occupation.
@@
Even if we assume the cross-occupation interaction terms $\beta_{2}$ and $\beta_{hw}$ are zero, the difference-in-differences estimator $\hat{\beta}_{h}$ from (8) identifies:
$\begin{split}\hat{\beta}_{h}=&\;\beta_{h}+\beta_{1}\cdot\Big[\underbrace{P(D^{o(i)}=1\mid D^{o(h)}=1,\text{Post})-P(D^{o(i)}=1\mid D^{o(h)}=1,\text{Pre})}_{\Delta P(D^{o(i)}=1\mid D^{o(h)}=1)}\\[2.5pt] &\hskip 31.0001pt-\underbrace{\big(P(D^{o(i)}=1\mid D^{o(h)}=0,\text{Post})-P(D^{o(i)}=1\mid D^{o(h)}=0,\text{Pre})\big)}_{\Delta P(D^{o(i)}=1\mid D^{o(h)}=0)}\Big]\\[5.0pt] &\;-\;\beta_{w}\cdot\Big(P(D^{o(i)}=1\mid D^{o(h)}=1,\text{Post})-P(D^{o(i)}=1\mid D^{o(h)}=0,\text{Post})\Big).\end{split}$ (10)
Even under the assumption that assortative matching is stable before and after COVID— i.e.,
$\Delta P(D^{o(i)}=1\mid D^{o(h)}=1)=\Delta P(D^{o(i)}=1\mid D^{o(h)}=0)=0,$
the estimator $\hat{\beta}_{h}$ will still load on the direct effect $\beta_{w}$ through differences in wives’ occupational exposure across treated and untreated husbands in the post period.
Section 6 presents several exercises designed to separate these two components and quantify the relative contributions of spillover and direct effects.